OpenStudy (anonymous):

evaluate the expression . sin 30 degrees + cos 60 degrees

OpenStudy (lgbasallote):

are you supposed to use trig formulas?

OpenStudy (anonymous):

Remember your 30 60 90 right triangle

OpenStudy (anonymous):

i dont know

jimthompson5910 (jim_thompson5910):

You can also use the unit circle http://www.math.ucsd.edu/~jarmel/math4c/Unit_Circle_Angles.png

jimthompson5910 (jim_thompson5910):

notice at the angle 30 degrees, the y coordinate is 1/2 so sin(30) = 1/2

OpenStudy (anonymous):

\[\text{Sin}[30{}^{\circ}]+\text{Cos}[60{}^{\circ}]=\frac{1}{2}+\frac{1}{2}=1 \]